find the roots of the polynomial equation:x^3-2x^2+10x+136=0

Answer by stanbon(75887)

find the roots of the polynomial equation:x^3-2x^2+10x+136=0
-----
---
I graph the cubic and found a zero at x = -4
----
Using synthetic division, find the other zeros:
-4)....1....-2....10....136
........1....-6....34...|..0
----
Quotient: x^2-6x+34
---
Find the zeros of the quotient:
x = [6 +- sqrt(6^2-4*34)]/2
---
x = [6 +- sqrt(36-136)]/2
---
x = [6 +- sqrt(-100)]/2
---
x = [2 +- 5i]
x = 2+5i or x = 2 - 5i or -4
===================================
Cheers,
Stan H.
===================